Two similar squares are in a ratio of 3:5. If a side of the larger square is 35 inches, what is the measure of a side of the smaller square?
Answer by lwsshak3(11628)

Two similar squares are in a ratio of 3:5. If a side of the larger square is 35 inches, what is the measure of a side of the smaller square?
**
let x=side of smaller square
..
ratio is that of areas:
3/5=x^2/35^2
3(35)^2=5x^2
3675=5x^2
x^2=3675/5=735
x=√735≈27.1 Inches